OpenStudy (anonymous):

Suppose the area of a circle is 16π. What is its circumference?

OpenStudy (acxbox22):

16pi=pi r^2 16=r^2 r=4

OpenStudy (anonymous):

midhum madhu answer my question you were wrong

OpenStudy (acxbox22):

c=2 x pi x r c= 2 x 4 x pi c=8 pi

OpenStudy (anonymous):

no midhum madhu tell me the correct answer

OpenStudy (midhun.madhu1987):